(Week 2 Problem 20.55 IP A capacitor has plates with an area of What is the potential difforeno between the plates when the charge on the apacor plates is 42d7 Will your answer to part A Decrease Request Answer Provide Feedback 으 Type here to search SA
Answer #1

Solution)

Given, Area of plate, A= 1.3*10^-2 m^2

Separation between plates, d= 0.81 * 10^-3 m

Dielectric constant, k= 1.8

Part a) Capacitance, C= kAEo/d = 1.8* (1.3*10^-2)*(8.85*10^-12)/(0.81*10^-3)

=2.43 *10^-10 F

Now, V= Q/C = 4.2*10^-6/2.43*10^-10 =17238.88 V (Ans)

=========

Part b)

Since, Potential difference V is inversely proportional to k

So, Potential difference decreases when dielectric constant increases.

============

Good luck!:)